# Filtering Pixel Latent Variables for Unmixing Volumetric Images

### 08 Dec 2023 v1

@@ Abstract
Measurements of different overlapping components require robust unmixing algorithms to convert the raw multi-dimensional measurements to useful unmixed images. Such algorithms perform reliable separation of the components when the raw signal is fully resolved and contains enough information to fit curves on the raw distributions. In experimental physics, measurements are often noisy, undersampled, or unresolved spatially or spectrally. We propose a novel method where bandpass filters are applied to the latent space of a multi-dimensional convolutional neural network to separate the overlapping signal components and extract each of their relative contributions. Simultaneously processing all dimensions with multi-dimensional convolution kernels empowers the network to combine the information from adjacent pixels and time- or spectral-bins, facilitating component separation in instances where individual pixels lack well-resolved information. We demonstrate the applicability of the method to real experimental physics problems using fluorescence lifetime microscopy and mode decomposition in optical fibers as test cases. The successful application of our approach to these two distinct experimental cases, characterized by different measured distributions, highlights the versatility of our approach in addressing a wide array of imaging tasks.
@@ 1 Introduction
Unmixing refers to the technique of breaking down a measurement into its unique constituents and their respective proportions. This concept is commonly linked to spectral analysis, where the overlapping spectral signatures of different emitters hinder the accurate identification of each emitter’s abundance, especially when the spectral signature is undetermined or the data is insufficiently sampled. Beyond spectral measurements, unmixing applies to any measurements that combine spatial dimensions (x,y,z) to additional continuous and often multi-channel measures. In fluorescence lifetime microscopy, this additional dimension is the time it took for the photon to be emitted[4]. In spectrally and spatially (S) resolved measurements for characterizing optical fibers, this dimension is the amplitude of the propagated light for different input wavelengths [21].
In this work, we propose a generalized unmixing method based on convolutional neural networks that applies to various types of multi-dimensional measurements (Fig. 1). We specifically focus on noisy data or undersampled distributions, for which common methods fail to accurately unmix the different contributions. The proposed method relies on transforming the overlapping input distributions to a latent space where they are untangled and can be separated using simple bandpass filters. We demonstrate the method on synthetic datasets based on real data and created for real experimental physics use cases, showcasing the general applicability of the approach.
2 Related work
In this work, we demonstrate an unmixing method on two experimental physics use cases: fluorescence lifetime microscopy (FLIM) and mode decomposition in optical fibers. In both cases, the input data is composed of overlapping distributions that need to be decomposed into single channel outputs to retrieve useful information.
2.1 Multidimensional CNNs and the latent space
@@
Three-dimensional Convolutional Neural Networks (3D CNNs) offer advantages over their two-dimensional counterparts in tasks involving volumetric imaging such as medical imaging [28] or spatiotemporal data such as videos [14]. Their ability to capture both spatial and temporal features simultaneously leads to improved representations of complex patterns in three-dimensional space [32]. Unmixing tasks face the challenge of combining 3D input with 2D multi-channels output, making the task difficult to adapt to existing 3D models; for example, with a 3D U-Net, where the 3D features learned in the encoder cannot be directly linked to the 2D features required in the decoder. CNNs that take as input a 3D volume and predict a 2D image is not a common architecture as it is a distinct task from volumetric segmentation, and no architecture has been developed specifically for this task. Tasks on 3D volumes are generally either in the volumetric space from input to output (3D segmentation [28]) or a classification layer is added to the learned representation for a 1D output (classification [27]). 2.5D networks have been proposed for the segmentation of volumes into 2D slice-wise segmentation for cases where 3D CNNs are not applicable because of slice discontinuities [23], or in cases where the data is too limited to train 3D networks [15]. However, 2.5D networks were shown to be less accurate than 3D CNNs when 3D CNNs are applicable [2].
In our method, the features are represented in the 3D space from the entry to the exit of the network, so that existing 3D architectures can be exploited for unmixing tasks. This leads to a general model that is not constrained in the number of target output channels, because the learned features are translated into channels directly from the learned 3D latent space. This latent space is an abstract and simplified representation of the data; in convolutional neural networks, the latent space simplifies the raw input pixel data into an informative representation that is used by the subsequent layers to perform the task[17].
2.2 Fluorescence lifetime microscopy
Fluorescence lifetime microscopy (FLIM) uses the arrival time of fluorescence photons to assign them to specific fluorophores or to measure environmental changes in biological samples [4]. The fluorescence decay follows an exponential probability distribution characterized by the parameter corresponding to the fluorescence lifetime [5]. The fluorescence lifetime represents a distinctive photophysical characteristic unique to each fluorescent molecule, which is used in FLIM to discriminate between several fluorophores. However, even by carefully selecting molecules based on their fluorescence lifetime, exponential distributions always overlap and signal unmixing can be challenging, especially in the low photon-count regime (100 photons/pixel) or when multiple fluorophores are present at a given position. For live-cell imaging, where photobleaching and sample movement should be minimized, FLIM measurements are generally associated with low photon-counts [31]. Additionally, measurements of interactions between multiple components increase the complexity of assigning each photon to a given emitting molecule.
2.2.1 FLIM analysis
Different unmixing approaches have been proposed for FLIM such as Maximum Likelihood Estimation (MLE) [3], phasor analysis [8], and deep learning based strategies [30]. MLE is an effective method for fitting parameters to describe a curve given a set of samples from the distribution and is a common approach for FLIM data analysis [3]. An alternative approach relying on phasor analysis has been proposed for unmixing fluorescence lifetime distributions [8]. The performance of both approaches is strongly limited by low photon counts (300) and when more than three components have to be unmixed [7, 6, 25, 8].
Recently, deep learning based approaches have been proposed for FLIM data analysis. The Fluorescence lifetime imaging network (FLI-Net) [30] is a 3D-CNN that takes as input a 3D FLIM volume (x,y,t) where each pixel is characterized by photons sampled in a biexponential distribution with aim to assign a mean lifetime to each pixel rather than unmixing the input image. The Fluorescence Lifetime AdderNet is a 1D network trained for lifetime unmixing [33], which can handle spatially distinct structures and does not use the spatial features to improve photon assignment in overlapping structures.
2.2.2 Unmixing strategies for fluorescence microscopy
A common method for unmixing multi-channel images from a single acquisition is to combine the lifetime and spectral information (spectral FLIM, S-FLIM). In this case, acquired volumes are 4D (x,y,,t). The multidimensional information facilitates the assignment of photons to a given fluorescent molecule, since emitters characterized by similar lifetimes can have different spectral signatures or vice versa. Phasor analysis has been explored as a method for unmixing S-FLIM data [19, 10]. Deep learning-based unmixing of S-FLIM data has also been demonstrated with UNMIX-ME [29], which uses convolutional neural networks to map noisy hyperspectral FLIM data to spatially independent unmixed fluorescence coefficient values. The method allows the separation of up to four components in simulated samples.
To our knowledge, there is a gap in the literature for FLIM unmixing method that applies to images in the low photon count regime (100 photon/pixel) and more than two fluorescent emitters per pixel. Such an approach would be transformative for the democratization of FLIM for multi-channel live-cell fluorescence microscopy and nanoscopy.
2.3 Mode decomposition in multi-mode fibers
Multi-mode fibers are optical fibers that support the propagation of multiple simultaneous light modes [11]. Spatially and spectrally resolved (S) imaging[21] is a measurement commonly used for characterizing how light travels through multi-mode optical fibers. By measuring the spatial distribution of the intensity at the output of the fiber for a range of laser wavelengths, the modes of propagation of the light through the fiber can be characterized. The 2D image at the fiber output is not merely a linear combination of each distinct mode because the coherent modes interfere constructively or destructively with each other in function of their relative phases. In a two modes fiber, the intensity of a single pixel versus the wavelength of the laser propagating through the fiber, within a small spectral range, oscillates periodically with a sinusoidal distribution (Fig. 7). When more than one mode propagates through the fiber, the intensity oscillates as the sum of each sinus function associated to each combination of two modes.
In a properly sampled measurement where the frequency of acquisition is above the Nyquist frequency, these oscillations allow to separate each mode with a Fourier transform based method [20] or by principal component analysis (PCA) [24, 26]. Other methods exist to achieve real-time mode decomposition from a single image with various algorithms [13, 1, 18], but they all require the a priori knowledge of the spatial distribution of each mode.
To our knowledge, no method for mode decomposition in multi-mode fibers solve both limitations of current approaches: separate modes with unknown spatial distributions, and successfully retrieve the contribution of each mode with a sub-Nyquist sampling rate.
@@ 3 Methodology
3.1 Datasets
Experiments are conducted with three datasets: 1) a modified version of the MNIST dataset [16] with overlaid digits and simulated lifetime (3D-MNIST), 2) a FLIM dataset generated from STimulated Emission Depletion (STED) microscopy images of neuronal proteins (FLIM-STED), and 3) a simulation of spectrally and spatially resolved measurements of light propagating in multi-mode fibers (S-MMF).
3D-MNIST The 3D-MNIST dataset was used to develop the method and evaluate the influence of different parameters. Images from the classes 4, 8 and 9 are overlaid to create mixed images. Random Poisson noise is generated over the whole image as a fourth target channel. This multi-dimensional version of MNIST is generated according to Algo. 1.
The generated MNIST dataset is parametrized by a signal-to-noise factor , decay constants for each channel (, , , ), a time dimension length , and a number of images . Each data sample is created by randomly sampling up to three digits from classes (4, 8, 9) from the MNIST training set. We chose the digits 4, 8 and 9 because this triplet of digits has the highest fraction of overlap (Supp. fig. 2). There is a probability 0.5 of picking one digit for each of the (4,8,9) classes (lines 9 to 13), such that a training sample can contain 0, 1, 2 or 3 digits. The images are independently randomly rotated, flipped and translated (line 17). To reduce the pixel counts from [0,255] to a noisier range more representative of real experimental conditions, the image is first normalized to the range [0,1] by dividing each image by its maximum value (line 19). A random count is assigned to each pixel by sampling an intensity value in a Poisson distribution parameterized by the pixel value, multiplied by the SNR factor (lines 21 and 22). For each count, an equivalent number of time values are sampled following an exponential distribution parameterized by the decay constant of the channel (line 24). The time values of each individual pixel are converted to a histogram of bins of fixed width (line 26). A background channel is also created by randomly sampling values in a Poisson distribution with parameter and creating a time histogram following the same method as for the digits channels (lines 29 to 33). The four volumes (4, 8, 9, background) are added along the channel dimension to create a mixed volume of dimensions () (line 35). The process is repeated times to generate the 3D-MNIST dataset.
The 3D-MNIST dataset counts 8,000 combined images generated with and , from which 6,400 randomly sampled images were used for training and 1,600 for testing.
STED-FLIM The STED-FLIM dataset consists of real microscopy images of four different proteins tagged with fluorescent markers in fixed cortical neurons (Suppl. sec. A). The signal from each fluorescent marker was spectrally separated to obtain ground truth images with four independent channels. The STED-FLIM dataset was generated by adapting the 3D-MNIST data generation algorithm (algo. 1). The following modifications were applied to the algorithm: 1) dataset is a set of 4-color FLIM images, 2) channels are not randomly discarded and independently sampled (lines 10-15), 3) the true measured photon counts were preserved instead of normalized and randomized (lines 16-22), 4) no background channel was added (lines 29-33). The lifetime values (Suppl. table 1) were chosen to fit within the range of commercially available fluorescent dyes for STED microscopy [12]. We chose values within this range so that the separation between lifetimes is variable (0.7 ns A-B, 0.6 ns B-C, 0.2 ns C-D). Since the lifetime distributions are synthetic, the actual lifetimes do not need to match a specific dye, so we attributed generic names (A, B, C and D) to the fluorescent markers for clarity. Channels A and C were acquired with confocal microscopy while channels B and D were acquired with STED microscopy. This allows to test if the method can adapt to varying photon counts and spatial resolutions. The STED-FLIM dataset counts 400 training and validation images which were preprocessed into 8,261 crops, and 20 test images.
| Protein | Marker | Lifetime [ns] |
| MAP2 | A | 2.6 |
| PSD95 | B | 3.3 |
| VGLUT1 | C | 3.9 |
| FUS | D | 4.1 |
S-MMF The S-MMF dataset consists in images of the intensity measured at the output of optical fibers for different input wavelengths [22]. The simulated measure has three dimensions; two spatial (2D images of the light distribution at the fiber output) and one spectral (each image represents the measure obtained from a different light wavelength). The dataset was computed with the equations described in Quiquempois et al. (2021) [24]. A vector finite difference mode-solver [9] was used to find mode properties around 1550 nm for various index profiles and fiber geometries. All fibers in the dataset support up to 6 modes. The mode spatial distribution and effective index dependency on wavelength (dispersion) were taken into account. For all data samples, we considered a spectral range of 6 nm for a fiber length of 5 m. The dataset includes measurements with various combinations of up to 6 propagating modes, four different fiber geometries (step-index with varying numerical apertures and core size, GRIN, W-type, H-type and a realistic step-index geometry with a central dip), and different mode-specific intensities (). We used 1500 fiber measurements to train the network and 175 for testing.
@@ 3.2 Model
Network architecture 3D convolution kernels allow the processing of all dimensions of a 3D input volume simultaneously. We chose to use a 3D CNN for unmixing undersampled data so that the 3D kernels can combine essential information from neighboring pixels and neighboring time- or spectral-bins. The network used here is a standard 3D U-Net (Supp. sec. B). As is standard for U-Nets, the dimensionality of the input remains unchanged through the layers; the input 3D volume results in a 3D output of the same size. We added a layer with one convolution filter at the end of the network that maps the output of the U-Net to a single-channel volumetric output of dimension , with and coming from the images spatial dimensions and being the pixel latent space size.
Latent space filters For unmixing applications, the input is a 3D volume and the target is a channel 2D image. In our method, called Latent Unmixing, we apply predefined bandpass filters to the pixel latent space of the network output to pool this dimension into channels. This forces the network to learn a mapping that transforms the input overlapping distributions into a latent space where the distributions do not overlap and can be separated with bandpass filters (Fig. 1).
@@
Algo. 2 describes the training process of the latent unmixing method. The training dataset must contain mixed volumetric data samples , along with unmixed ground truth images for each of ground truth channels. At each epoch, the latent representation of a first batch of data samples is generated with the 3D U-Net model (line 6). Predefined bandpass filters are applied to this latent vector to aggregate the output values included within each filter to the corresponding channel (lines 8 and 9). The resulting multi-channel prediction is normalized along the channel dimension so that each pixel is represented by its probability of belonging to this channel (line 11). The loss is the mean square error between the predicted () and ground truth () multi-channel maps of probabilities. The loss is backpropagated to update the weights of the network using stochastic gradient descent. This process is repeated over all batches of training data, and over a number of epochs .
4 Experiments and results
4.1 MNIST with simulated decay
@@
We first apply the Latent Unmixing approach to untangle the signal from overlapping 4, 8, 9 digits and background in the 3D-MNIST dataset. In the images, the SNR (parameter in Algo. 1) is such that a simple 2D U-Net can not accurately separate the digits from the intensity input only (Supp. sec. D), demonstrating that the time dimension is needed to assign individual photons to the right channel. The decay parameters were chosen as 0.1, 0.2, 0.3 and 0.4 for digits 4, 8, 9 and the background respectively (Fig. 2, top-left). Our Latent unmixing method learns to map the input distributions sampled in the decay function to a pixel latent space where the contributions are well separated to fit the predefined bandpass filers (Fig. 2 and 3). The latent space is filtered with 4 bandpass filters of width 7 equally spaced along the pixel latent vector of length 28. The method can separate the digits accurately, achieving a Pearson correlation coefficient between true and predicted pixel intensities in the range above 0.82 for all channels over the 1,600 images from the test set (Fig. 4). To compare with curve fitting methods, we used MLE to unmix the same test set pixel-by-pixel. We fix the decay constants as the real values, and use MLE to fit the relative abundance. MLE achieved Pearson correlation coefficients between -0.1062 and 0.2484, showing no or negligible correlation between true and predicted intensities (Supp. sec. E and Table 2). This result was expected, since MLE is not an accurate method for separating as many as 4 colors with pixel counts in the range [0,20]; it is better suited for unmixing less colors with much higher pixel counts [3].

### 06 Apr 2024 v2

@@ Abstract
The development of robust signal unmixing algorithms is essential for leveraging multimodal datasets acquired through a wide array of scientific imaging technologies, including hyperspectral or time-resolved acquisitions. In experimental physics, enhancing the spatio-temporal resolution or expanding the number of detection channels often leads to diminished sampling rate and signal-to-noise ratio, significantly affecting the efficacy of signal unmixing algorithms. We propose applying band-pass filters to the latent space of a multi-dimensional convolutional neural network to disentangle overlapping signal components, enabling the isolation and quantification of their individual contributions. Using multi-dimensional convolution kernels to process all dimensions simultaneously enhances the network’s ability to extract information from adjacent pixels, time- or spectral-bins. This approach enables more effective separation of components in cases where individual pixels do not provide clear, well-resolved information. We showcase the method’s practical use in experimental physics through two test cases that highlight the versatility of our approach: fluorescence lifetime microscopy and mode decomposition in optical fibers. The latent unmixing method extracts valuable information from complex signals that cannot be resolved by standard methods. Application of latent unmixing to real FLIM experiments will increase the number of distinguishable fluorescent markers. It will also open new possibilities in optics and photonics for multichannel separations at increased sampling rate.
@@ 1 Introduction
Signal unmixing refers to the technique of breaking down a measurement into its unique constituents and their respective proportions. This concept is often associated with spectral analysis, where the challenge lies in accurately identifying the abundance of each emitter from their overlapping spectral signatures. Beyond spectral data, signal unmixing can apply to any measurements that rely on the combination of multiple dimensions or components. The difficulty to accurately separate the components becomes more pronounced when the specific spectrum of an emitter is unknown or when the data coverage is insufficient.
In this work, we propose an unmixing method that applies to various types of underdefined multi-dimensional measurements that outperforms classical curve-fitting based methods such as Maximum Likelihood Estimation[4]. Our method, called Latent Unmixing, transforms the overlapping input contributions to a latent space where they are untangled and can be separated by applying filters directly in the latent space (Fig. 1).
1.1 Related work
Linear unmixing methods exploit the fact that the signal measured at each pixel is a linear combination of each components’ specific signature weighted by their relative abundance. When the signature is known, the relative abundance parameters can be retrieved using estimation algorithms such as least squares fitting methods[41, 19] and Maximum Likelihood Estimation (MLE) [4]; given that the signal is sufficiently sampled for the curves to be accurately estimated from the raw data. Another linear unmixing approach, the Phasor method[24], translates the raw measurements to a 2D phasor space where the distributions are represented by the real and imaginary components of the Fourier transform. This representation then allows to estimate the relative abundance of different components through a linear interpolation, knowing the phasor coordinates of individual components obtained through control measurements.
Deep learning methods have been proposed for the task of unmixing images, especially for hyperspectral data[17, 8, 50]. Hyperspectral data has a high spectral sampling obtained at the expense of diminished spatial resolution[46]. Although this limitation poses its own challenges, mainly that single pixels cover large ground areas and therefore often contain several materials [51, 14], the methods developed do not cover the inverse problem of measurements that are well resolved spatially but undersampled spectrally. The deep learning models for processing hyperspectral data – either for classification[1, 53, 23], super-resolution[26], denoising[40], or unmixing[47, 34] – are generally based on three-dimensional Convolutional Neural Networks (3D CNNs) to process the raw hyperspectral measurements as volumetric data.
@@ 1.1 Related work
3D CNNs offer advantages over their two-dimensional counterparts in tasks involving volumetric datasets [45] or spatiotemporal acquisitions [22]. Their ability to capture and combine spatial and temporal features simultaneously leads to improved representations of complex patterns in three-dimensional space [49]. For signal unmixing of 3D imaging data, it is required that the CNN takes as input a 3D volume and predicts the proportion of n component in each pixel, resulting in a stack of n 2D images. Signal unmixing tasks present the significantly different challenge of converting a 3D input (e.g. (x, y, t)) into a 2D multi-channel output (e.g. n (x, y)), which can not directly be solved with existing end-to-end 3D models, such as the 3D U-Net. Tasks on 3D volumes are generally either in the volumetric space from input to output (3D segmentation [45]) or a classification layer is added to the learned representation for a 1D output (classification [33]).
We designed our method over the intuition that a 3D U-Net can be optimized to map the input data from the input space, where the signal components are undersampled and overlapping, to a space where the different contributions are distinctly differentiated. Other works previously used the distangled representation of input data in the latent space of an autoencoder to unmix hyperspectral data[38, 32]. Those methods are however specific to hyperspectral data, because they either require precise reference spectra[38] and therefore do not account for possible spectral variability[9], or process a single pixel at a time and do no use the spatial and spectral information from neighbouring pixels[32]. For hyperspectral data, the pixel spatial range is sufficiently large that there may be no need to rely on information from neighboring pixels to infer the components present within the pixel.
1.2 Contributions
The differentiating contributions of this work are as follows: 1) we developed a method to map input volumetric measurements to a latent space where the included signals are disentangled, 2) we specifically focus on the less studied unmixing problem for data that is well-resolved spatially but undersampled in its third dimension and outperform commonly used methods for low-sampling regimes, 3) we showcase the applicability of the method over three simulated datasets for two completely distinct experimental use cases and include a demonstration on real acquired data, which is enabled by the generalizability of the model to different parametrization of the input distributions.
@@ 2.1 Model
We chose to use a 3D U-Net for unmixing undersampled data so that the 3D kernels can combine essential information from neighboring pixels and neighboring time- or spectral-bins, since 3D convolution kernels allow the processing of all dimensions of a 3D input volume simultaneously. The network used here is a 3D U-Net with 4 downsampling blocks and 4 upsampling blocks (Suppl. sec. B). To generalize the model to contexts where the input distribution parameters vary, we developed a conditional paradigm where the parameters that define the distribution of the input volume is embedded to the encoder blocks. The whole architecture is described in details in Supp. sec. B. As is standard for U-Nets, the dimensionality of the input remains unchanged through the layers; the input 3D volume results in a 3D output of the same size. We added a layer with one convolution filter at the end of the network that maps the output of the U-Net to a single-channel volumetric output of dimension , with and coming from the images spatial dimensions and being the pixel latent space size. The receptive field of the final layer spans 110 pixels in width, height, and depth, contrasting with pixelwise approaches for which the predictive capacity is restricted by the very limited information of a single pixel.
2.2 Latent space filters
In convolutional neural networks, the latent space simplifies the raw input pixel data into an informative representation that is used by the subsequent layers to perform the task[25]. Our latent unmixing method uses this abstract and simplified representation of the data in the latent space to translate the 3D input into a multi-channel output.
@@ 2.2 Latent space filters
In our method, called Latent Unmixing, we apply predefined bandpass filters to the pixel latent space of the network output to pool this dimension into channels. This forces the network to learn a mapping that transforms the input overlapping distributions into a latent space where they do not overlap and can be separated with bandpass filters (Fig. 1). Algo. 1 describes the training process of the Latent Unmixing method. The training dataset must contain mixed volumetric data samples , along with unmixed ground truth images for each of ground truth channels. At each epoch, the latent representation of a first batch of data samples is generated with the 3D U-Net model (line 6). Predefined bandpass filters are applied to this latent vector to aggregate the output values included within each filter to the corresponding channel (lines 8 and 9). The resulting multi-channel prediction is normalized along the channel dimension so that each pixel is represented by its probability of belonging to this channel (line 11). The loss is the mean square error between the predicted () and ground truth () multi-channel maps of probabilities. The loss is backpropagated to update the weights of the network using stochastic gradient descent. This process is repeated over all batches of training data, and over a number of epochs .
3 Experiments
In this work, we demonstrate a signal unmixing method on a variation of the MNIST dataset and on two experimental physics use cases: fluorescence lifetime microscopy (FLIM) and mode decomposition in optical fibers. In all three cases, the input data is composed of noisy, undersampled, and overlapping distributions that need to be decomposed into multi-channel outputs.
3.1 3D MNIST with simulated decay
Dataset The 3D-MNIST dataset was used to benchmark the method’s performance. Images from the digit classes 4, 8 and 9 are overlaid to create mixed images. We chose the digits 4, 8 and 9 because this triplet of digits has the highest fraction of overlap (Supp. fig. 2). Random Poisson noise is generated over the whole image as a fourth target channel to simulate background noise in optical microscopy. A multi-dimensional version of the mixed images of the MNIST digits is generated according to Algo. 2. The generated dataset is parametrized by a signal-to-noise factor , a decay constants for each channel (, , , ), a time dimension histogram of bins, and a number of images . Each mixed image is created by randomly sampling up to three digits from classes (4, 8, 9) from the original MNIST training set. There is a probability 0.5 of picking one digit for each of the (4,8,9) classes (lines 9 to 13), such that a training sample can contain 0, 1, 2 or 3 digits. The digits are independently randomly rotated, flipped and translated (line 17). The mixed images are first normalized to the range [0,1] by dividing each image by its maximum value (line 19). To more accurately mirror real experimental conditions, normalized pixel values are adjusted by sampling a random count for each pixel from a Poisson distribution, with the sampling parameter being the pixel value multiplied by the signal-to-noise ratio (SNR) factor (lines 21 and 22). For each count, a corresponding time is sampled following an exponential distribution parameterized by the decay constant of the channel (line 24). The time values of each individual pixel are converted to a histogram of bins of fixed width (line 26). A background channel is also created by randomly sampling values in a Poisson distribution with parameter and creating a time histogram following the same method as for the digits channels (lines 29 to 33). The four volumes (4, 8, 9, background) are added along the channel dimension to create a mixed volume of dimensions () (line 35). The process is repeated times to generate the 3D-MNIST dataset. The 3D-MNIST dataset counts 8,000 mixed volumes generated with and , from which 6,400 randomly sampled images were used for training and 1,600 for testing.
3.1.1 Results
@@
We first apply the Latent Unmixing approach to untangle the signal from overlapping 4, 8, 9 digits and background in the 3D-MNIST dataset. In the images, the SNR (parameter in Algo. 2) is such that a simple 2D U-Net can not accurately separate the digits from the intensity input only (Supp. sec. D), demonstrating that the time dimension is needed to assign individual photons to the right channel. The decay parameters were chosen as 0.1, 0.2, 0.3 and 0.4 for digits 4, 8, 9 and the background respectively (Fig. 2, top-left). Our Latent Unmixing method learns to map the input distributions sampled in the decay function to a pixel latent space where the contributions are well separated to fit the predefined bandpass filters (Fig. 2). Four bandpass filters of width 7 equally spaced along the pixel latent vector of length 28 are used. The method can separate the digits accurately, achieving a Pearson correlation coefficient between true and predicted pixel intensities in above 0.82 for all channels over the test set (Fig. LABEL:fig:mnistb). We compared the result of Latent Unmixing with the established curve fitting approach MLE to unmix the same test set pixel-by-pixel. We fix the decay constants as the actual values, and use MLE to fit the relative abundance. MLE achieved Pearson correlation coefficients between -0.1062 and 0.2484, showing no or negligible correlation between true and predicted intensities (Supp. sec. E and Table 1). This result was expected, since MLE is not an accurate method for separating as many as 4 channels with limited pixel counts in the range [0,20] [4].